### v1

@@ 0.C.2 Solver Performance on Individual Benchmarks
Alt. Loops REI.id491.ep67924 SAT 0.6 6.6 SAT 0.2 12.8
Alt. Loops REI.id491.ep70134 SAT 0.6 7.0 SAT 0.4 23.2
Alt. Loops REI.id491.ep96326 SAT 0.7 6.8 SAT 5.6 39.0
Alt. Loops REI.id502.ep71027 SAT 0.6 6.9 SAT 56.7 42.9
Alt. Loops REI.id502.ep71445 SAT 0.6 6.9 SAT 15.9 46.3
Alt. Loops REI.id502.ep84337 SAT 0.5 6.7 SAT 2.1 30.2
Alt. Loops REI.id502.ep86449 SAT 0.9 6.8 SAT 21.3 42.0
Alt. Loops REI.id502.ep90122 SAT 0.8 6.9 SAT 266.6 45.8
Alt. Loops REI.id512.ep70217 SAT 0.4 6.9 SAT 1.3 41.2
Alt. Loops REI.id512.ep70283 SAT 0.7 6.7 SAT 1.0 33.3
Alt. Loops REI.id512.ep75804 SAT 0.7 7.1 SAT 986.8 50.4
Alt. Loops REI.id512.ep97330 SAT 0.7 7.0 SAT 0.1 11.2
Alt. Loops REI.id512.ep97385 SAT 0.7 6.7 SAT 567.2 47.5
Alt. Loops REI.id518.ep68431 UNS 101.2 9.2 UNS 523.8 46.8
Alt. Loops REI.id518.ep70223 UNS 80.0 9.4 UNS 1358.8 49.7
Alt. Loops REI.id518.ep95639 UNS 126.9 10.7 UNS 132.8 39.7
Alt. Loops REI.id518.ep95900 UNS 112.7 10.2 UNS 602.4 45.1
Alt. Loops REI.id518.ep96115 SAT 83.8 10.2 UNS 561.7 45.2

### v2

@@ 0.C.2 Solver Performance on Individual Benchmarks
Alt. Loops REI.id491.ep67924 SAT 0.4 6.6 SAT 0.2 12.8
Alt. Loops REI.id491.ep70134 SAT 0.5 6.5 SAT 0.4 23.2
Alt. Loops REI.id491.ep96326 SAT 0.4 6.1 SAT 5.6 39.0
Alt. Loops REI.id502.ep71027 SAT 0.4 6.5 SAT 56.7 42.9
Alt. Loops REI.id502.ep71445 SAT 0.4 6.5 SAT 15.9 46.3
Alt. Loops REI.id502.ep84337 SAT 0.6 7.0 SAT 2.1 30.2
Alt. Loops REI.id502.ep86449 SAT 0.8 6.9 SAT 21.3 42.0
Alt. Loops REI.id502.ep90122 SAT 0.4 6.4 SAT 266.6 45.8
Alt. Loops REI.id512.ep70217 SAT 0.6 6.6 SAT 1.3 41.2
Alt. Loops REI.id512.ep70283 SAT 0.6 6.8 SAT 1.0 33.3
Alt. Loops REI.id512.ep75804 SAT 0.5 6.4 SAT 986.8 50.4
Alt. Loops REI.id512.ep97330 SAT 0.6 6.8 SAT 0.1 11.2
Alt. Loops REI.id512.ep97385 SAT 0.4 6.3 SAT 567.2 47.5
Alt. Loops REI.id518.ep68431 UNSAT 97.2 9.9 UNSAT 523.8 46.8
Alt. Loops REI.id518.ep70223 UNSAT 85.6 9.2 UNSAT 1358.8 49.7
Alt. Loops REI.id518.ep95639 UNSAT 125.9 10.2 UNSAT 132.8 39.7
Alt. Loops REI.id518.ep95900 UNSAT 123.8 9.9 UNSAT 602.4 45.1
Alt. Loops REI.id518.ep96115 UNSAT 89.9 9.9 UNSAT 561.7 45.2